# Decentralized Small Gain and Phase Stability Conditions for Grid-Forming Converters: Limitations and Extensions

### v1

@@ I Introduction
To overcome these limitations, recent research has proposed decentralized stability criteria [2, 3, 4] that enable the impact of a single converter to be assessed independently of others in the system. These conditions are designed such that, if each individual converter satisfies them, the overall system stability is guaranteed without requiring explicit analysis of converter-converter interactions. This paradigm shift offers a scalable and modular pathway to stability certification in converter-dominated grids, reducing reliance on exhaustive system-wide simulations. Moreover, these decentralized conditions yield valuable insights for converter control design to ensure system stability.
A well-known decentralized criterion for ensuring stability requires that all system components exhibit passivity. This requirement is starting to appear in upcoming grid codes for grid-forming converters [5]. Although this condition is naturally satisfied by the network, it does not hold for converters. At high frequencies, passivity can be enforced through suitably control strategies. At low frequencies, understood here in the synchronous reference (dq) frame, however, non-passivity is intrinsically associated with the constant-power characteristics of converters, a property that cannot be removed.
@@
Another relevant study [9] formulates the system in a polar reference frame,in which GFM converters are passive at low frequencies. Assuming time-scale separation, the authors propose a hybrid criterion: passivity is verified in the polar frame at low frequencies, while the standard rectangular admittance representation is used at high frequencies. However, the time-scale separation assumption may not hold in practice or may be difficult to verify. Decentralized parametric conditions on P-f and Q-V control were also proposed in [4]. However, the scope is limited to predefined control structure and ignore lower level converter dynamics.
@@
Data and code for reproducing the results are available at [10].
@@ II Mixed Gain-Phase Stability Conditions
A matrix is said to be sectorial if $W(A)$ lies entirely within an angular sector of the complex plane and $0\notin W(A)$. For a sectorial matrix $A$, there exists an invertible matrix $T$ such that $A=T^{*}DT$, where $D$ is a unique diagonal matrix (up to permutation).
The phases of $A$ are defined as the arguments of the diagonal entries of $D$, ordered so that $\overline{\phi}(A)$ and $\underline{\phi}(A)$ denote the maximum and minimum phases, with $\overline{\phi}(A)-\underline{\phi}(A)<\pi$. The phase interval $\underline{\phi}(A),\,\overline{\phi}(A)]$ is shorty denoted by $\Phi(A)$. This definition can be extended to quasi-sectorial (i.e. $0\in\partial W(A)$) and to semi-sectorial (i.e. $0\in\partial W(A)$ and $\overline{\phi}(A)-\underline{\phi}(A)\leq\pi$) matrices [13]. Geometrically, these phases correspond to the angular boundaries of the numerical range. In direct analogy with singular values, maximum and minimum phases can be used to bound the phases of the eigenvalues of a matrix product as in (3).
@@ II-B Decentralized condition for power system stability
In the context of power systems, we are interested in assessing the stability of the interconnection of $n$ power electronic converters to a given network. Let
@@
    $\max_{i}\overline{\phi}(Y_{C_{i}}(j\omega))<\pi-\overline{\phi}(\mathbf{Y}_{net}^{-1}(j\omega)),$
    $\min_{i}\underline{\phi}(Y_{C_{i}}(j\omega))>-\pi-\underline{\phi}(\mathbf{Y}_{net}^{-1}(j\omega))$
@@ IV-A Power-Polar coordinates frame
With these definitions, the converter’s new input–output relation becomes as in 12 and similar the transformation is applied to the network, yielding 13.
@@
This change of variables can be interpreted as a loop transformation, as illustrated in Fig. 4. Assuming the transformation is well-defined, input/output stability of the transformed system and the original system are equivalent. Indeed, the stability is determined by the zeros of $\det(Y_{net}(s)+Y_{C}(s))$. Since $\det(\mathbf{J}_{C}(s)+\mathbf{J}_{net})=\det(\mathcal{E}(Y_{net}(s)+Y_{C}(s))\mathcal{F})$, the zeros of the transformed system and the original are identical. Moreover, if there are no right plane pole/zero cancellations, then input–output stability implies internal stability, as shown for instance in [4]. Theorem II.2 can be applied directly to the transformed system without loss of generality.
@@
In this frame, the converter admittance at zero frequency is quasi-sectorial with a well-defined phase. While this may not extend across the low-frequency range, it remains a promising indication. Nevertheless, for the small-phase criterion to be applicable, sectoriality must hold across the entire frequency range for both $J_{C}(s)$, and $J^{-1}_{net}(s)$. While the transformation can resolve the low-frequency non-sectoriality, it does not guarantee that sectoriality will persist at higher frequencies.
@@ IV-B Unified Approach
Ideally, at low frequencies the matrices should correspond to those in (10)-(11) (denoted from now on as $\mathcal{E}_{J}$, $\mathcal{F}_{J}$ and $\mathcal{C}_{J}$), while at high frequencies they should revert to the rectangular frame (i.e., $\mathcal{E}=\mathcal{F}=I$ and $\mathcal{C}=0$, denoted from now on as $\mathcal{E}_{Y}$, $\mathcal{F}_{Y}$ and $\mathcal{C}_{Y}$). A continuous transition between these two regimes is necessary. A trivial choice for $\mathcal{E}(s)$, and similar for $\mathcal{F}(s)$ and $\mathcal{C}(s)$, can be as in (15) using an low/high-pass filtering approach. The filters transfer function are given as: $H_{LPF}(s)=w_{c}/(s+w_{c})$ and $H_{HPF}(s)=1-H_{LPF}(s)$, with $w_{c}$ the filter cut-off frequency.
@@
This transformation proves in general inadequate since sectoriality near the filter cut-off frequency $w_{c}$.
To ensure sectoriality while obtaining tight phase bounds, we propose instead the transformation in (17), with an additional degree of freedom $W_{i}$. In this formulation, frequency-dependent filtering is applied exclusively to $\mathcal{C}$ and $\mathcal{F}(s)$, while $\mathcal{E}(s)$ remains constant.
@@ V-A Virtual Admittance compensation
A tighter bound can be obtained by reducing the large phase spread of the numerical range, and therefore limiting the extreme combinations of phases. Since the converter implements a virtual admittance control $Y_{v}(s)$, and its behavior around $50$ Hz is mainly determined by the virtual admittance, we can post-multiply $Y_{C}(s)$ by the inverse of $Y_{v}(s)$, such that $Y_{C}(s)Y_{v}(s)^{-1}\approx I$ around those frequencies. The network admittance is transformed accordingly to $Y_{net}(s)Y_{v}(s)^{-1}$. From another viewpoint, a known dynamic behaviour is shifted from the converter to the network, thereby avoiding the conservatism. Figure 6 illustrates the effect of this compensation. It can be seen that applying the proposed virtual admittance filtering leads to noticeably improved bounds. This figure also highlights the necessity of the translation term $\mathcal{C}$. Since one eigenvalue’s phase approaches $\pi$ at low frequencies and pre-/post-multiplication does not alter it, we must rely on the small gain at those frequencies to ensure stability. Conversely, with a properly chosen $\mathcal{C}$, the stability assessment can rely solely on the phase criterion, as discussed in the next subsection.
At first glance, this approach may appear to undermine the decentralized nature of the certificate. However, in practice, exact knowledge of the virtual admittance values (that is, the behaviour around 5050Hz) is not required, and approximate values are sufficient, as shown afterwards. In a decentralized setting, these values can either be selected a priori as fixed parameters or assigned differently across units by the stability certifier. In any case, proposed grid codes for grid-forming converters already define fixed ranges for the effective impedance [17].
@@ V-B Full transformation
In Figure 7 the small-phase criteria is shown with the frequency-depended transformation (15), which matrices are defined as in (17). Compared to the previous results, we can obtain sectoriality for the full frequency range and the closed-loop system stability is guaranteed by the small-phase conditions.
@@ V-C Unstable case
To further validate the method, we examine the sufficient condition of the theorem: if the closed-loop system is unstable, then the small-phase and the small-gain condition must be simultaneously violated at some frequency. In practice, destabilizing a grid-forming converter is not straightforward. Following [18], we introduce a purely resistive load at the PCC and increase the infinite bus impedance. By switching the control mode to constant-Q and reducing the SVM damping of the GFM from $0.5$ to $0.05$, the closed-loop becomes unstable, exhibiting sustained oscillations at 1.2 Hz. Figure 8 illustrates the mixed gain/phase test, considering the proposed transformation. Enabling the reactive power control reduces the gain at low frequencies, leading to the small-gain condition to be satisfied. However, while in the stable case ($D=0.5$) the small-phase condition is satisfied, in the unstable case ($D=0.05$), the converter phase increases substantially, and together with the network phase causes the phase condition to be violated, while the gain condition is also not satisfied around the unstable oscillation frequency.
@@ VI Example 2 - IEEE 14 Bus System
Figure 8: GFM Infinite Bus - Stable vs Unstable.
To extend and show the scalability of our approach, we consider the modified IEEE 14 Bus System [10]. Four GFM converters are connected at the buses 2, 3, 6, and 8. Similarly to the infinite bus example, to build an unstable case, we select a low damping value for converter 8, connect purely resistive loads, and increase the line 8-9 impedance. In this setup, the system is unstable with an unstable oscillation at 0.54 Hz. In Figure 9, the decentralized conditions are shown. While the converters 2, 3, and 6 satisfy the small-phase condition, converter 8 does not. In particular, below 1.64 Hz, its phase overlaps the network phase. This does not imply the instability of the system; however, knowing that the system is unstable, it identifies the responsible converter, and by appropriately retuning its control parameters, stability can potentially be restored. For example, if we apply the same tuning of converter 6, which satisfy the small-phase condition, to converter 8, then stability is recovered.
In this example, each converter implements a different virtual admittance value, while the same value is applied to all converters for the transformation in (17). This shows that, even when the exact virtual admittance is unknown, it is still possible to use a reasonable estimate to obtain adequate phase bounds and thus an effective stability validation with the proposed approach.
Figure 9: IEEE14 Bus system
We intentionally excluded GFL converters. Although the concepts presented are general and applicable to different control structures, GFL converters exhibit distinct dynamics that require a separate, more detailed analysis and potentially a different definition of the transformation matrices. Furthermore, as highlighted in [3], the small-gain condition plays a more critical role, making it necessary to shift the high gain of GFMs toward the network to enhance its strength.
@@ VII Conclusion
The method was demonstrated on a grid-forming converter connected to an infinite bus, where virtual admittance filtering further reduced conservativeness, and on the IEEE 14-bus system, where it successfully identified unstable devices. While the results are promising, challenges remain, particularly when studying systems that include both grid-forming and grid-following converters, as well as in guaranteeing stability and sectoriality of the grid impedance. Future work will focus on systematic selection of transformation matrices and on reducing conservativeness by embedding network topology information into the method.
@@ Appendix A GFM Control structure
In this paper, we consider the GFM control structure represented in Figure 10. The controller is made in the stationary reference frame $\alpha\beta$. It consists of a Proportional-Resonant (PR) current controller and a Virtual Admittance ($Y_{v}$) voltage controller. On top of these, the power control loop is implemented through a Virtual Synchronous Machine and a Proportional-Integral (PI) reactive power controller. Unless explicitly stated, the reactive power controller is disabled, and a constant voltage magnitude reference is used. Parameters and further details are available in [10].

### v2

@@ Abstract
[…]
This work has been funded by the CETPartnership, the Clean Energy Transition Partnership under the 2023 joint call for research proposals, co-funded by the European Commission (GA N 101069750) and with the funding organizations AEI (National State Research Agency, Spain), PtJ (MWIKE) (Projekträger Jülich/Forschungszentrum Jülich GmbH, Germany), GSRI (General Secretariat for Research and Innovation, Greece), SEAI (Sustainable Energy Authority of Ireland), FFG (Austrian Research Promotion Agency) and RCN (The Research Council of Norway).
@@ I Introduction
To overcome these limitations, recent research has proposed decentralized stability criteria [3, 14, häberle2025decentralizedparametricstabilitycertificates] that enable the impact of a single converter to be assessed independently of others in the system. These conditions are designed such that, if each individual converter satisfies them, the overall system stability is guaranteed without requiring explicit analysis of converter-converter interactions. This paradigm shift offers a scalable and modular pathway to stability certification in converter-dominated grids, reducing reliance on exhaustive system-wide simulations. Moreover, these decentralized conditions yield valuable insights for converter control design to ensure system stability.
A well-known decentralized criterion for ensuring stability requires that all system components exhibit passivity. This requirement is starting to appear in upcoming grid codes for grid-forming converters [12]. Although this condition is naturally satisfied by the network, it does not hold for converters. At high frequencies, passivity can be enforced through suitably control strategies. At low frequencies, understood here in the synchronous reference (dq) frame, non-passivity is intrinsically associated with the constant-power characteristics of converters, a property that cannot be removed.
@@
Another relevant study [10] formulates the system in a polar reference frame, in which grid-forming (GFM) converters are passive at low frequencies. Assuming time-scale separation, the authors propose a hybrid criterion: passivity is verified in the polar frame at low frequencies, while the standard rectangular admittance representation is used at high frequencies. However, the time-scale separation assumption may not hold in practice or may be difficult to verify. Decentralized parametric conditions on P-f and Q-V control were also proposed in [häberle2025decentralizedparametricstabilitycertificates]. However, the scope is limited to predefined control structure and ignore lower level converter dynamics.
@@
The remainder of this paper is organized as follows: Section II reviews decentralized gain-phase stability criteria, while Section III addresses their limitations regarding GFM converters. Section IV details our proposed extensions. The methodology is validated in Sections V and VI for a converter connected to an infinite bus and the IEEE 14 Bus system, respectively. Finally, Section VII concludes the paper.
Data and code to reproduce the results are available at [7].
@@ II Mixed Gain-Phase Stability Conditions
A matrix is said to be sectorial if $W(A)$ lies entirely within an angular sector of the complex plane and $0\notin W(A)$. For a sectorial matrix $A$, there exists an invertible matrix $T$ such that $A=T^{*}DT$, where $D$ is a unique diagonal matrix (up to permutation). The phases of $A$ are defined as the arguments of the diagonal entries of $D$, ordered so that $\overline{\phi}(A)$ and $\underline{\phi}(A)$ denote the maximum and minimum phases, with $\overline{\phi}(A)-\underline{\phi}(A)<\pi$. The phase interval $\underline{\phi}(A),\,\overline{\phi}(A)]$ is shortly denoted by $\Phi(A)$. This definition can be extended to quasi-sectorial (i.e. $0\in\partial W(A)$) and to semi-sectorial (i.e. $0\in\partial W(A)$ and $\overline{\phi}(A)-\underline{\phi}(A)\leq\pi$) matrices [6]. Geometrically, these phases correspond to the angular boundaries of the numerical range. In direct analogy with singular values, maximum and minimum phases can be used to bound the phases of the eigenvalues of a matrix product as in (3).
@@ II-B Decentralized condition for power system stability
Figure 2: Illustration of a multi-converter power system
In the context of power systems, we are interested in assessing the stability of the interconnection of $n$ power electronic converters to a given network, illustrated in Figure 2. Let
@@
    $\overline{\phi}(Y_{C_{i}}(j\omega))<\pi-\overline{\phi}(\mathbf{Y}_{net}^{-1}(j\omega))\quad\forall i$
    $\underline{\phi}(Y_{C_{i}}(j\omega))>-\pi-\underline{\phi}(\mathbf{Y}_{net}^{-1}(j\omega))\quad\forall i$
    $\max_{i}\overline{\phi}(Y_{C_{i}}(j\omega))-\min_{i}\underline{\phi}(Y_{C_{i}}(j\omega))<\pi$
@@ IV-A Power-Polar coordinates frame
With these definitions, the converter’s new input–output relation becomes as in (12) and similar the transformation is applied to the network, yielding (13).
@@
[…]
@@
In this frame, the converter admittance at zero frequency is quasi-sectorial with a well-defined phase. While this may not extend across the low-frequency range, it remains a promising indication. Nevertheless, for the small-phase criterion to be applicable, sectoriality must hold across the entire frequency range for both $\mathbf{J}_{C}(s)$, and $\mathbf{J}^{-1}_{net}(s)$. While the transformation can resolve the low-frequency non-sectoriality, it does not guarantee that sectoriality will persist at higher frequencies.
@@ IV-B Unified Approach
Ideally, at low frequencies the matrices should correspond to those in (10)-(11) (denoted from now on as $\mathcal{E}_{J,i}$, $\mathcal{F}_{J,i}$ and $\mathcal{C}_{J,i}$), while at high frequencies they should revert to the rectangular frame (i.e., $\mathcal{E}=\mathcal{F}=I$ and $\mathcal{C}=0$, denoted from now on as $\mathcal{E}_{Y,i}$, $\mathcal{F}_{Y,i}$ and $\mathcal{C}_{Y,i}$). A continuous transition between these two regimes is necessary. A trivial choice for $\mathcal{E}_{i}(s)$, and similar for $\mathcal{F}_{i}(s)$ and $\mathcal{C}_{i}(s)$, can be as in (16) using an low/high-pass filtering approach. The filters transfer function are given as: $H_{LPF}(s)=w_{c}/(s+w_{c})$ and $H_{HPF}(s)=1-H_{LPF}(s)$, with $w_{c}$ the filter cut-off frequency.
@@
This transformation proves in general inadequate since sectoriality is lost near the filter cut-off frequency $w_{c}$. To ensure sectoriality while obtaining tight phase bounds, we propose instead the transformation in (17), with an additional degree of freedom $W_{i}$. In this formulation, frequency-dependent filtering is applied exclusively to $\mathcal{C}$ and $\mathcal{F}(s)$, while $\mathcal{E}(s)$ remains constant.
@@ V-A Virtual Admittance compensation
A tighter bound can be obtained by reducing the large phase spread of the numerical range, and therefore limiting the extreme combinations of phases. Since the converter implements a virtual admittance control $Y_{v}(s)$, and its behavior around $50$ Hz is mainly determined by the virtual admittance, we can post-multiply $Y_{C}(s)$ by the inverse of $Y_{v}(s)$, such that $Y_{C}(s)Y_{v}(s)^{-1}\approx I$ around those frequencies. The network admittance is transformed accordingly to $Y_{net}(s)Y_{v}(s)^{-1}$. From another viewpoint, a known dynamic behaviour is shifted from the converter to the network, thereby avoiding the conservatism. In (17), this correspond in the the selection $W_{i}(s)=Y_{v}(s)^{-1}$.
Figure 7 illustrates the effect of this compensation. It can be seen that applying the proposed virtual admittance filtering leads to noticeably improved bounds.
At first glance, this approach may appear to undermine the decentralized nature of the certificate. However, in practice, exact knowledge of the virtual admittance values (that is, the behavior around 5050Hz) is not required, and approximate values are sufficient, as shown afterwards. In a decentralized setting, these values can either be selected apriori as fixed parameters or assigned differently across units by the stability certifier. In any case, proposed grid codes for grid-forming converters already define fixed ranges for the effective impedance [11].
Remark 4.
Figure 7 shows one eigenvalue approaches $\pi$ as the frequency tends to zero. In this region, the small-gain condition does not ensure $|\lambda|<1$, rendering stability assessment inconclusive. This motivates the shift introduced by $\mathcal{C}$, which relocates the eigenvalues of the open-loop transfer function in a more suitable region where small-phase allows to certify stability.
@@ V-B Full transformation
In Figure 8, the small-phase criteria is shown with the frequency-depended transformation (15), with matrices defined as in (17). The cut-off frequency of the filters is set to $0.5$ Hz. Compared to the previous results, we can obtain sectoriality for the full frequency range and the closed-loop system stability is guaranteed by the small-phase conditions alone. This means that the stability of this GFM converter does not depend on the grid strength, meant as its gain, and stability is preserved under both weak and strong conditions. In this sense, the small-phase condition provides a grid-strength-independent stability guarantee. However, variations in the network characteristics, such as changes in the $R/X$ ratio or loading, may alter the network phase. Such variations can reduce the phase margin and potentially drive the converter toward instability.
@@ V-C Unstable case
To further validate the method, we examine the sufficient condition of the theorem: if the closed-loop system is unstable, then the small-phase and the small-gain condition must be simultaneously violated at some frequency. In practice, destabilizing a grid-forming converter is not straightforward. Following [8], we introduce a purely resistive load at the PCC and increase the infinite bus impedance. By switching the control mode to constant-Q and reducing the VSM damping of the GFM from $0.5$ to $0.05$, the closed-loop becomes unstable, exhibiting sustained oscillations at 1.2 Hz. Figure 9 illustrates the mixed gain/phase test, considering the proposed transformation. Enabling the reactive power control reduces the gain at low frequencies, leading to the small-gain condition to be satisfied. However, while in the stable case ($D=0.5$) the small-phase condition is satisfied, in the unstable case ($D=0.05$), the converter phase increases substantially, and together with the network phase causes the phase condition to be violated, while the gain condition is also not satisfied around the unstable oscillation frequency.
Figure 9: GFM Infinite Bus - Stable ($D=0.5$) vs Unstable ($D=0.05$).
@@ VI Example 2 - IEEE 14 Bus System
Figure 10: IEEE 14 Bus System diagram
To extend and show the scalability of our approach, we consider the modified IEEE 14 Bus System shown in Figure 10, which parameters are given in [7]. Four GFM converters are connected at the buses 2, 3, 6, and 8. Similarly to the infinite bus example, to build an unstable case, we select a low damping value for converter 8, connect purely resistive loads, and increase the line 7-8 impedance. In this setup, the system is unstable with an unstable oscillation at 0.54 Hz driven by converter 8. In Figure 11, the decentralized conditions are shown using the proposed transformation and virtual admittance compensation. While the converters 2, 3, and 6 satisfy the small-phase condition, converter 8 does not. In particular, below 1.64 Hz, its phase overlaps the network phase. This does not imply the instability of the system; however, knowing that the system is unstable, it identifies the responsible converter, and by appropriately retuning its control parameters, stability can potentially be restored. For example, if we apply the same tuning of converter 6, which satisfy the small-phase condition, to converter 8, then stability is recovered.
In this example, each converter implements a different $R/X$ ratio in its virtual admittance, as reported in Table I, while a fixed value ($R_{v}=0.01$, $X_{v}=0.1$) is used for the virtual admittance compensation for all converters. This demonstrates that, even when the exact virtual admittance is unknown, a reasonable estimate can still be used to obtain adequate phase bounds and, consequently, an effective stability validation with the proposed approach.
Figure 11: Stability analysis of the 14 Bus System
TABLE I: Virtual admittance setting in 14 Bus System
GFM Bus $R_{v}$ (p.u) $X_{v}$ (p.u)
22 $0.02$ $0.1$
33 $0.05$ $0.1$
66 $0.03$ $0.1$
88 $0.01$ $0.1$
@@ VII Conclusion
The method was demonstrated on a grid-forming converter connected to an infinite bus, where virtual admittance filtering further reduced conservativeness, and on the IEEE 14-bus system, where it successfully identified unstable devices.
While the results are promising, challenges remain, particularly when studying systems that include not only grid-forming converters but also other devices, such as synchronous machines and grid-following converters. Synchronous machines have similar dynamics to the GFM converter considered here and we expect the proposed approach to work. However, the impact of auxiliary control loops (e.g. exciters, automatic voltage regulators, power system stabilizers) requires further verification. For grid-following converters, which exhibit distinct dynamics, the small-gain condition plays a more critical role, and the benefits of the proposed transformation are limited. To obtain a less conservative gain condition for GFL units, it would be beneficial to shift the high gain of the GFMs toward the network, thereby enhancing its strength.
Future work will focus on systematic selection of transformation matrices and on reducing conservativeness by embedding network topology information into the method.
@@ Appendix A GFM Control structure
In this paper, we consider the GFM control structure represented in Figure 12. The controller are implemented in the stationary reference frame $\alpha\beta$. It consists of a Proportional-Resonant $PR(s)$ current controller, given in (18), and a Virtual Admittance $Y_{v}(s)$ voltage controller, given in (19), with virtual resistance $R_{v}$ and virtual inductance $L_{v}$.
$PR(s)=K_{p}+\frac{K_{i}}{s^{2}+(2\pi 50)^{2}}$ (18)
$Y_{v}(s)=\frac{1}{sL_{v}+R_{v}},\quad L_{v}=\frac{X_{v}}{2\pi 50}$ (19)
The synchronization control is implemented through a Virtual Synchronous Machine (VSM) with inertia $H$ and damping $D$. A Proportional-Integral (PI) controller regulates the reactive power changing the reference voltage magnitude. Unless explicitly stated, the reactive power controller is disabled, and a constant voltage magnitude reference is used. Parameters and further details are available in [7].